# MLKV: Multi-Layer Key-Value Heads for Memory Efficient Transformer Decoding

### v1

@@ Abstract
Zayd Muhammad Kawakibi Zuhri1                        Muhammad Farid Adilazuarda2
Ayu Purwarianti1                        Alham Fikri Aji2
1Institut Teknologi Bandung (ITB), 2MBZUAI Correspondence: zaydzuhri@gmail.com
@@ 1 Introduction
The most notable methods for handling large KV caches are the approaches that directly reduce the number of KV heads used, from here on referred to as KV sharing. Multi-Query Attention (MQA) by Shazeer (2019) uses only a single key and value projection for all attention heads in a layer. It reduces memory bandwidth for KV cache by $1/n\_heads$, which is significant, but results in some degradation in quality and stability. Grouped-Query Attention (GQA) from Ainslie et al. (2023) addresses these issues and introduce an intermediate solution of sharing a number KV heads to multiple groups of attention heads. This effectively reduces KV cache size by $n\_groups/n\_heads$. They also introduce an uptraining scheme to convert any transformer checkpoint to MQA or GQA. With a reasonable number of heads, GQA can achieve near parity on benchmarks with a vanilla model. However, the reduction in KV cache size is still limited to $1/n\_heads$ with MQA, which still might not be enough for some applications.
@@ 2.3 Grouped-Query Attention (GQA)
Then Ainslie et al. (2023) proposed to generalize the KV sharing to groups of query heads. This allows for a more flexible configuration and a better performance trade-off. Given $g$ groups of query heads each having one KV head, GQA is defined as follows with $g<h$:
@@ 2.4 KV Caching
The reason why these KV sharing methods were developed was to reduce memory overhead of the KV cache. KV caching is used to optimize autoregressive inference by only keeping the key and value activations of previous tokens. This way, only the singular newest token needs to be passed into the model to generate the next token in the sequence. Given batch size $b$ and sequence length $s$, the KV cache has the dimensions $[2,b,s,l,h,d_{k}]$. Thus, it scales linearly with batch size, sequence length, and model size, which means it can grow indefinitely during inference. Taking OPT-175B as an example, its parameters require 325GB of memory. Yet when inferencing a sequence length of 2048 and computing a batch of 128 generations at once, the KV cache can take up to 950GB of memory (Liu et al., 2023).
@@ 3 Multi-Layer Key-Value (MLKV)
Method KV Cache Size (# Elements)
MHA $2bslhd_{k}$
MQA $2bsld_{k}$
GQA $2bslgd_{k}$
MLKV $2bsmgd_{k}$
Table 1: Theoretical KV cache sizes by number of elements for each KV sharing method. $b$ for batch size, $s$ for sequence length, $l$ for number of layers, $m$ for number of layers with their own KV heads, $h$ for number of (query) heads, $g$ for number of KV head groups, and $d_{k}$ for head dimension. Note that $m<l$ and $g<h$.
@@ 4.5 Evaluation Method
Aside from benchmark performance, inference time metrics must also be measured for each model. Specifically, we evaluate memory usage and throughput. This is done by a custom script that loads and runs the models up to their limits and measures the metrics at the same time. A dummy KV cache with length 2000 is initialized, then the model is autoregressively run token by token for 48 tokens, up to the sequence length limit. This process is timed to obtain tokens/second. Right after exiting the forward pass, memory usage is measured via the NVIDIA System Management Interface (SMI). The output is the VRAM usage measured in megabytes. We deduct the background and model memory usage from this reading. Do note that due to this, the measurement results might seem to go out-of-memory prematurely, because the memory taken for tensor computations are not considered.

### v2

@@ Abstract
Zayd Muhammad Kawakibi Zuhri1, Muhammad Farid Adilazuarda2, Ayu Purwarianti1, Alham Fikri Aji2 1Institut Teknologi Bandung  2MBZUAI zaydzuhri@gmail.com
@@ 1 Introduction
The most notable methods for handling large KV caches are the approaches that directly reduce the number of KV heads used, from here on referred to as KV sharing. Multi-Query Attention (MQA) by Shazeer (2019) uses only a single key and value projection for all attention heads in a layer. It reduces memory bandwidth for KV cache by $1/n\_heads$, which is significant, but results in some degradation in quality and stability. Grouped-Query Attention (GQA) from Ainslie et al. (2023) addresses these issues and introduces an intermediate solution of sharing several KV heads to multiple groups of attention heads. This effectively reduces KV cache size by $n\_groups/n\_heads$. They also introduce an uptraining scheme to convert any transformer checkpoint to MQA or GQA. With a reasonable number of heads, GQA can achieve near parity on benchmarks with a vanilla model. However, the reduction in KV cache size is still limited to $1/n\_heads$ with MQA, which still might not be enough for some applications.
@@ 2.3 Grouped-Query Attention (GQA)
Ainslie et al. (2023) then proposed to generalize the KV sharing to groups of query heads. This allows for a more flexible configuration and a better performance trade-off. Given $g$ groups of query heads each having one KV head, GQA is defined as follows with $g<h$:
@@ 2.4 KV Caching
The reason why these KV sharing methods were developed was to reduce the memory overhead of the KV cache. KV caching is used to optimize autoregressive inference by only keeping the key and value activations of previous tokens. This way, only the singular newest token needs to be passed into the model to generate the next token in the sequence. Given batch size $b$ and sequence length $s$, the KV cache has the dimensions $[2,b,s,l,h,d_{k}]$. Thus, it scales linearly with batch size, sequence length, and model size, which means it can grow indefinitely during inference. Taking OPT-175B (Zhang et al., 2022) as an example, its parameters require 325GB of memory. Yet when inferencing a sequence length of 2048 and computing a batch of 128 generations at once, the KV cache can take up to 950GB of memory (Liu et al., 2023).
@@ 3 Multi-Layer Key-Value (MLKV)
Method KV Cache Size (# Elements) Cache Size of OPT-175B (GB)
MHA $2bslhd_{k}$ 144.0
MQA $2bsld_{k}$ 1.5
GQA $2bslgd_{k}$ 36.0
MLKV $2bsmgd_{k}$ 0.375
Table 1: Theoretical KV cache sizes by number of elements for each KV sharing method. $b$ for batch size, $s$ for sequence length, $l$ for number of layers, $m$ for number of layers with their own KV heads, $h$ for number of (query) heads, $g$ for number of KV head groups, and $d_{k}$ for head dimension. Note that $m<l$ and $g<h$. Following the example in 2.4, column 3 shows calculated cache size of OPT-175B if KV shared, given $b=8$, $s=1024$, $g=24$ for GQA and $g=1$, $m=24$ for MLKV, in float16 precision.
@@ 4.5 Evaluation Method
Aside from benchmark performance, inference time metrics are also measured for each model. Specifically, we evaluate memory usage and throughput. This is done by a script that loads and runs the models up to their limits and measures the metrics at the same time. A dummy KV cache with length 2000 is initialized, then the model is autoregressively run token by token for 48 tokens, up to the sequence length limit. This process is timed to obtain tokens/second. Right after exiting the forward pass, memory usage is measured via the NVIDIA System Management Interface (SMI). The output is the VRAM usage measured in megabytes. We deduct the background and model memory usage from this reading. Note that due to this, the measurement results might seem to go out-of-memory prematurely, because the memory taken for tensor computations are not considered.
@@ 6 Discussion
7 Related Work
KV cache optimization methods on dimensions other than $h$ and $l$ have also been done. Approaches in reducing the sequence length $s$ of the cache do it by compressing information in the context. Attention mask variations like sliding windows (Zaheer et al., 2020), dilated sliding windows (Beltagy et al., 2020), and attention sinks (Xiao et al., 2023) attempt to limit the receptive field of tokens to some length smaller than the actual sequence length, with the assumption that the information of the tokens before will be compressed to some other positions. A more concrete compression-based solution is SCISSORHANDS (Liu et al., 2023) which only keeps pivotal tokens with the assumption of persistence of importance, such that the KV cache length can be kept short. FastGen (Ge et al., 2024) employs an adaptive KV cache compression strategy based on specific policies on special tokens, punctuation, locality, and frequency of different positions to determine which KV heads to prune.
Moreover, we are aware of the contemporaneous work on Cross-Layer Attention (CLA) (Brandon et al., 2024), published when we were nearing the end of finishing our work on MLKV. CLA was experimented on larger models trained from scratch, instead of the uptraining practice in our work, which utilizes available model checkpoints. Additional ablations were done with CLA, including non-uniform layer sharing patterns, but missing more extreme configurations like MLKV-2 and MLKV-1. The ablations are also based on a KV cache memory budget equivalence, instead of matching parameter counts in MLKV.